# A Comprehensive Characterization of Galaxy-Cool CGM Connections at $z<0.4$ with DESI Year 1 Data

### v1

@@ 3.1 Dependencies on galaxy properties
With the large number of galaxy-quasar pairs across a wide range of galaxy properties, we first investigate the Ca II rest equivalent width as a function of impact parameter and stellar mass, comparing star-forming and quiescent galaxies. The results are shown in Figure 3 with the left and right panels indicating the Ca II rest equivalent width as a function of impact parameter around star-forming and quiescent galaxies, respectively, with different stellar masses indicated by the colors. The number of pairs is summarized in Table 3.1. We find several trends which are described in the following:
@@
$r_{p}$ Star-forming $0.08_{-0.06}^{+0.32}$ $\phantom{0}8.1_{-1.6}^{+0.4}$ $\phantom{0}0.61\pm 1.23$ $-2.15\pm 0.20$ 139946
$\phantom{0}9.0_{-0.5}^{+0.5}$ $\phantom{0}1.76\pm 0.46$ 263731
$\phantom{0}9.9_{-0.4}^{+0.6}$ $\phantom{0}6.79\pm 0.57$ 285305
$10.7_{-0.2}^{+0.8}$ $12.43\pm 1.81$ 47469
$0.06_{-0.04}^{+0.09}$ $\phantom{0}8.1_{-1.6}^{+0.4}$ $\phantom{0}0.28\pm 1.30$ 139874
$\phantom{0}9.0_{-0.5}^{+0.5}$ $\phantom{0}1.78\pm 0.52$ 249580
$\phantom{0}9.8_{-0.3}^{+0.7}$ $\phantom{0}5.70\pm 0.89$ 147123
$10.7_{-0.2}^{+0.8}$ $12.94\pm 5.09$ 12413
$0.21_{-0.06}^{+0.19}$ $\phantom{0}9.5_{-1.0}^{+0.2}$ $\phantom{0}4.93\pm 1.13$ 32145
$10.1_{-0.4}^{+0.4}$ $\phantom{0}8.25\pm 0.95$ 120188
$10.7_{-0.2}^{+0.8}$ $12.49\pm 2.02$ 35056
Quiescent $0.15_{-0.13}^{+0.25}$ $\phantom{0}9.6_{-1.6}^{+0.4}$ $\phantom{0}6.62\pm 2.85$ $-1.86\pm 0.32$ 61976
$10.4_{-0.4}^{+0.3}$ $\phantom{0}4.66\pm 1.65$ 134889
$10.9_{-0.2}^{+1.0}$ $\phantom{0}8.01\pm 2.25$ 116110
$0.09_{-0.07}^{+0.08}$ $10.0_{-2.0}^{+0.5}$ $\phantom{0}4.34\pm 2.83$ 117276
$10.8_{-0.3}^{+1.1}$ $\phantom{0}8.79\pm 3.67$ 61702
$0.24_{-0.07}^{+0.16}$ $10.3_{-1.7}^{+0.2}$ $\phantom{0}6.31\pm 2.78$ 31112
$10.9_{-0.4}^{+1.1}$ $\phantom{0}5.50\pm 1.57$ 102885
AGN $0.07_{-0.05}^{+0.33}$ $10.2_{-3.2}^{+1.3}$ $\phantom{0}4.80\pm 1.27$ $-2.49\pm 1.29$ 39559
$r_{p}/r_{\mathrm{vir}}$ Star-forming $0.20_{-0.18}^{+0.20}$ $\phantom{0}8.8_{-2.3}^{+0.4}$ $\phantom{0}4.05\pm 0.93$ $-2.29\pm 0.24$ 69724
@@
$10.6_{-0.4}^{+0.9}$ $\phantom{0}6.21\pm 0.74$ 143798
Quiescent $0.22_{-0.20}^{+0.18}$ $\phantom{0}9.7_{-1.7}^{+0.3}$ $\phantom{0}7.25\pm 4.48$ $-1.90\pm 0.47$ 28399
$10.5_{-0.5}^{+0.2}$ $\phantom{0}5.13\pm 1.29$ 143912
$11.0_{-0.3}^{+1.0}$ $\phantom{0}1.08\pm 0.62$
@@
where $A_{100}$ is the equivalent width around galaxies at $r_{p}=100\,\mathrm{kpc}$. We first perform a global fitting to obtain the best-fit $\alpha$ parameter for each type of galaxies, assuming that the radial variations for the same type of galaxies are the same. We then fix the $\alpha$ parameters and obtain the best-fit amplitude for each stellar mass bin. Note that we do not include measurements within 10 kpc as the sight-lines are expected to also intersect the ISM of foreground galaxies, affecting our interpretations for the CGM. The best-fit profiles are shown by the color solid lines in Figure 3 with the best-fit parameter values listed in Table 3.1. The best-fit amplitudes, $A_{100}$, as a function of stellar mass are shown in Figure 4 with the colors representing the corresponding subsamples as in Figure 3. The results demonstrate that Ca II absorption strengths around star-forming galaxies at $\sim 100\,\mathrm{kpc}$ exhibit a mass dependence which can be described by
@@
Galaxy type dependence: In addition to the difference of mass dependence, for massive galaxies ($>10^{10}M_{\odot}$) with similar mass, at $r_{p}<30\rm\,kpc$, the absorption strengths around star-forming galaxies are higher than those around quiescent galaxies. By way of illustration, in the right panel of Figure 3, we show the Ca II measurements star-forming galaxies with $10^{10.7}M_{\odot}$ and $10^{9.9}M_{\odot}$. Comparing with the two highest stellar-mass bins of quiescent galaxies, the Ca II absorption within 50 kpc around $10^{10.7}M_{\odot}$ and $10^{9.9}M_{\odot}$ star-forming galaxies is twice as strong as the absorption around $10^{10.9}M_{\odot}$ and $10^{10.4}M_{\odot}$ quiescent galaxies, despite the star-forming galaxies having lower masses. This indicates that the cool gas distributes around star-forming and quiescent galaxies differently when one controls for stellar mass.
SFR dependence: We further examine the dependence of Ca II absorptions on star formation activity by dividing star-forming galaxies into low and high SFR groups using the equation: $\log_{10}{(\mathrm{SFR}/M_{\odot}\,\mathrm{yr^{-1}})}=0.5+0.95\,\left[\log_{10}{(M_{*}/M_{\odot})}-10.0\right].$ The samples for both SFR groups are then divided into three mass bins with matched median masses of $10^{9}$, $10^{9.9}$, and $10^{10.6}\,M_{\odot}$. We follow the previous steps to obtain the Ca II absorption strength as a function of impact parameter, fitting the trend globally with $\alpha=-2.15$. The fitted $A_{100}$ values are shown in Figure 5. By simultaneously fitting $A_{100}$ and SFR across the three mass bins, we find a positive correlation between Ca II absorption strength and SFR, following
$A_{100}=D(M_{*}/M_{\odot})\times\mathrm{SFR}^{0.27\pm 0.03},$ (4)
with $D(10^{9})=2.27\pm 0.24\,\mathrm{m\AA}$, $D(10^{9.9})=5.47\pm 0.25\,\mathrm{m\AA}$ and $D(10^{10.6})=7.72\pm 0.62\,\mathrm{m\AA}$. This result further supports the connection between the cool CGM and the star formation activity of host galaxies.
@@
[…]
@@
Figure 8: Best-fit amplitudes as a function of median stellar masses for two redshift bins. The solid and open points represent low- and high-redshift samples respectively. The solid and dashed lines in the left panel are the best-fit power laws for the low- and high-redshift samples.
Figure 9: Dependence of Ca II on the azimuthal angles for galaxies with $\log_{10}(M_{*}/M_{\odot})\geq 9.5$. Examples of the the galaxy-quasar pairs are shown at the top: from left to right are edge-on galaxies categorized into three azimuthal angles: disk, between disk and pole, pole; followed by face-on galaxies. The galaxies are located at the centre, while the quasars are indicated by the cyan arrows. In the bottom panel, the grey band shows the measurements of the face-on samples.
@@ 3.2 Redshift dependence
With the correlations between Ca II absorption and galaxy properties explored, we now investigate the correlation between Ca II absorption and redshift. We separate star-forming and quiescent galaxies at $z=0.15$ and $z=0.17$, respectively, and measure Ca II absorption as a function of stellar mass. The left and right panels of Figure 7 show the results for star-forming and quiescent galaxies, while the upper and lower panels show the results for the low- and high-redshift bins respectively. We also obtain the best-fit amplitudes of power laws (Eq. 2) with fixed $\alpha$ parameters obtained previously for star-forming and quiescent galaxies. The best-fit amplitudes are shown in Figure 8. The solid and open data points indicate the measurements at low and high redshifts respectively.
@@
$A_{100}=C(z)\left(\frac{M_{*}}{10^{10}\,M_{\odot}}\right)^{\beta},$ (5)
and find $\beta=0.45\pm 0.06$. The best-fit values of $C$ are $6.05\pm 0.67\ \mathrm{m\AA}$ ($z<0.15$) and $7.07\pm 0.52\ \mathrm{m\AA}$ ($z>0.15$). This results in
$C(z)=(5.69\pm 0.89)\,\mathrm{m\AA}\,(1+z)^{1.11\pm 0.95}.$ (6)
We note that despite the uncertainty, the redshift trend is similar to the redshift evolution observed in the cool gas around galaxies traced by Mg II absorption lines (e.g., Lan, 2020; Schroetter et al., 2021). For example, Lan (2020) showed that the covering fraction of Mg II absorption lines with $W_{\lambda 2796}>1\,\rm\AA$ around star-forming galaxies evolve with redshift as $(1+z)^{2.2\pm 0.4}$. On the other hand, we do not observe any redshift dependence for Ca II around quiescent galaxies which is inconsistent with the results of Lan (2020) for quiescent galaxies. This might indicate that the cool gas around passive and star-forming galaxies evolve differently. However, further confirmation of such a scenario requires higher S/N measurements of Ca II absorption.
In Lan (2020), the author proposed that the redshift evolution of Mg II covering fraction is connected with the redshift evolution of star-formation rate of galaxies (e.g., Whitaker et al., 2012; Scoville et al., 2017; Koprowski et al., 2024) given that they have similar redshift evolution. For our Ca II measurements, we also estimate the SFR of star-forming galaxies at two redshift bins with stellar mass $>10^{10}M_{\odot}$ and find that the ratio between the median SFR, $\mathrm{SFR}(z>0.15)/\mathrm{SFR}(z<0.15)$, is $\sim 1.8$. Together with our results indicating a SFR and absorption strength correlation with $\propto\rm SFR^{0.27}$ (Fig. 5), the absorption strength difference ratio is expected to be $\sim 1.8^{0.27}\sim 1.17$, which is consistent with the redshift trend observed in Figure 8. This consistency suggests that the redshift evolution of Ca II absorption around star-forming galaxies might be closely related to the star-formation activities of galaxies.
@@ 3.3 Azimuthal angle dependence
One can further investigate possible mechanisms contributing to gas observed in the CGM by measuring gas absorption around star-forming galaxies as a function of azimuthal angle. This is motivated by the fact that galaxy simulations predict that galactic outflows driven by supernova and black-hole feedback preferentially escape galaxies along the minor axis (e.g., Nelson et al., 2019). To do so, we examine the azimuthal dependence of our sample by choosing star-forming galaxies with $\log_{10}\left({M_{*}/M_{\odot}}\right)\geq 9.5$ and with inclination angle larger than $40^{\circ}$. The inclination angles of galaxies and azimuthal angles between galaxy-quasar pairs are calculated based on the shape information measured via Tractor [^4] https://www.legacysurvey.org/dr9/catalogs/ (Lang et al., 2016) applied to the DESI Legacy Imaging Surveys (Dey et al., 2019). In general, the Tractor performs model-based photometry by simultaneously fitting sources across multiple bands and classifies them as point or extended sources with the point spread function taken into account. Extended sources are further modeled using exponential, de Vaucouleurs, or composite profiles to best represent their light distribution. We then divide the selected sample into three azimuthal angle bins: $|\phi|\in[0^{\circ},30^{\circ}]$, $[30^{\circ},60^{\circ}]$ and $[60^{\circ},90^{\circ}]$, which correspond to the disk, between disk and pole, and the pole of the foreground galaxies. Galaxy image examples can be found in the upper panel of Figure 9, where the foreground galaxies are located at the centers, while the background quasars are pointed out by cyan arrows.
@@
The results are shown in the lower panel of Figure 9 with colors indicating measurements for different azimuthal angles and the grey band for the measurements of the face-on galaxies. While there is a mild trend showing that the absorption along the major axis decreases faster than the absorption along the minor axis toward larger impact parameters, there is no significant correlation between Ca II absorption and the azimuthal angle. This result is inconsistent with previous results of the cool CGM obtained at higher redshifts, showing preferentially stronger Mg II absorption along the minor axis than along the major axis within 50 kpc (e.g., Bordoloi et al., 2011; Lan et al., 2014; Lan & Mo, 2018). In addition, our results are inconsistent with Ca II measurements from Zhu & Ménard (2013b). The authors reported that the absorption strength of Ca II is stronger along the minor axis than the major one. However, we note that the conclusion in Zhu & Ménard (2013b) is driven by a single measurement at $r_{p}\sim 15\ \mathrm{kpc}$, a trend that is not observed in our measurements with more pairs. We will further discuss the implications of our azimuthal angle measurements in Section 4.2.

### v2

@@ 3.1 Dependencies on Galaxy Properties
With the large number of galaxy–quasar pairs across a wide range of galaxy properties, we first investigate the Ca II rest equivalent width as a function of impact parameter and stellar mass, comparing star-forming and quiescent galaxies. The results are shown in Figure 3 with the left and right panels indicating the Ca II rest equivalent width as a function of impact parameter around star-forming and quiescent galaxies, respectively, with different stellar masses indicated by the colors. The number of pairs is summarized in Table 1. We find several trends, which are described in the following:
@@
$r_{p}$ Star-forming $0.08_{-0.06}^{+0.32}$ $\phantom{0}8.1_{-1.6}^{+0.4}$ $\phantom{0}0.61\pm 1.23$ $2.15\pm 0.20$ 139,946
$\phantom{0}9.0_{-0.5}^{+0.5}$ $\phantom{0}1.76\pm 0.46$ 263,731
$\phantom{0}9.9_{-0.4}^{+0.6}$ $\phantom{0}6.79\pm 0.57$ 285,305
$10.7_{-0.2}^{+0.8}$ $12.43\pm 1.81$ 047,469
$0.07_{-0.05}^{+0.08}$ $\phantom{0}8.9_{-0.6}^{+0.5}$ $\phantom{0}1.50\pm 0.47$ 261,176
$\phantom{0}9.8_{-0.4}^{+0.6}$ $\phantom{0}4.88\pm 0.78$ 166,954
$10.6_{-0.2}^{+0.6}$ $13.85\pm 3.74$ 018,175
$0.22_{-0.07}^{+0.18}$ $10.0_{-0.6}^{+0.4}$ $\phantom{0}6.33\pm 2.60$ 130,437
$10.6_{-0.2}^{+0.6}$ $\phantom{0}9.43\pm 6.73$ 048,104
Quiescent $0.15_{-0.13}^{+0.25}$ $\phantom{0}9.6_{-1.6}^{+0.4}$ $\phantom{0}6.62\pm 2.85$ $1.86\pm 0.32$ 061,976
$10.4_{-0.4}^{+0.3}$ $\phantom{0}4.66\pm 1.65$ 134,889
$10.9_{-0.2}^{+1.0}$ $\phantom{0}8.01\pm 2.25$ 116,110
$0.10_{-0.08}^{+0.07}$ $10.3_{-0.8}^{+0.7}$ $\phantom{0}7.94\pm 0.82$ 142,711
$1.1_{-0.1}^{+0.6} 11$$1.68\pm 1.60$ 013,032
$0.25_{-0.08}^{+0.15}$ $10.7_{-0.5}^{+0.3}$ $\phantom{0}4.63\pm 1.83$ 091,412
$11.2_{-0.2}^{+0.5}$ $14.88\pm 3.64$ 034,505
AGN $0.07_{-0.05}^{+0.33}$ $10.2_{-3.2}^{+1.3}$ $\phantom{0}4.80\pm 1.27$ $2.49\pm 1.29$ 039,559
$r_{p}/r_{\mathrm{vir}}$ Star-forming $0.20_{-0.18}^{+0.20}$ $\phantom{0}8.8_{-2.3}^{+0.4}$ $\phantom{0}4.05\pm 0.93$ $2.29\pm 0.24$ 069,724
@@
$10.6_{-0.4}^{+0.9}$ $\phantom{0}6.21\pm 0.74$ 143,798
Quiescent $0.22_{-0.20}^{+0.18}$ $\phantom{0}9.7_{-1.7}^{+0.3}$ $\phantom{0}7.25\pm 4.48$ $1.90\pm 0.47$ 028,399
$10.5_{-0.5}^{+0.2}$ $\phantom{0}5.13\pm 1.29$ 143,912
$11.0_{-0.3}^{+1.0}$ $\phantom{0}1.08\pm 0.62$ 364,133
@@
where $A_{100}$ is the equivalent width around galaxies at $r_{p}=100\,\mathrm{kpc}$. We first perform a global fitting to obtain the best-fit $\alpha$ parameter for each type of galaxy, assuming that the radial variations for the same type of galaxies are the same. We then fix the $\alpha$ parameters and obtain the best-fit amplitude for each stellar mass bin. Note that we do not include measurements within 10 kpc as the sightlines are expected to also intersect the ISM of foreground galaxies, affecting our interpretations for the CGM. The best-fit profiles are shown by the color solid lines in Figure 3 with the best-fit parameter values listed in Table 1. The best-fit amplitudes, $A_{100}$, as a function of stellar mass are shown in Figure 4 with the colors representing the corresponding subsamples as in Figure 3. The results demonstrate that Ca II absorption strengths around star-forming galaxies at $\sim 100\,\mathrm{kpc}$ exhibit a mass dependence that can be described by
@@
Galaxy type dependence: In addition to the difference of mass dependence, for massive galaxies ($10^{10}M_{\odot}$) with similar mass, at $r_{p}<30\rm\,kpc$, the absorption strengths around star-forming galaxies are higher than those around quiescent galaxies. By way of illustration, in the right panel of Figure 3, we show the Ca II measurement star-forming galaxies with $10^{10.7}M_{\odot}$ and $10^{9.9}M_{\odot}$. Comparing with the two highest stellar mass bins of quiescent galaxies, the Ca II absorption within 50 kpc around $10^{10.7}M_{\odot}$ and $10^{9.9}M_{\odot}$ star-forming galaxies is twice as strong as the absorption around $10^{10.9}M_{\odot}$ and $10^{10.4}M_{\odot}$ quiescent galaxies, despite the star-forming galaxies having lower masses. This indicates that the cool gas distributes around star-forming and quiescent galaxies differently when one controls for stellar mass.
SFR dependence: We further examine the dependence of Ca II absorptions on star formation activity by dividing star-forming galaxies into low- and high- SFR groups using the equation: $\log_{10}{(\mathrm{SFR}/M_{\odot}\,\mathrm{yr^{-1}})}=0.5+0.95\,\left[\log_{10}{(M_{*}/M_{\odot})}-10.0\right].$ The samples for both SFR groups are then divided into three mass bins with matched median masses of $10^{9}$, $10^{9.9}$, and $10^{10.6}\,M_{\odot}$. We follow the previous steps to obtain the Ca II absorption strength as a function of impact parameter, fitting the trend globally with $\alpha=-2.15$. The fitted $A_{100}$ values are shown in Figure 5. By simultaneously fitting $A_{100}$ and SFR across the three mass bins, we find a positive correlation between Ca II absorption strength and SFR, following
$A_{100}=C\left(\frac{M_{*}}{10^{10}\,M_{\odot}}\right)^{\beta}\left(\frac{\mathrm{SFR}}{M_{\odot}\mathrm{yr^{-1}}}\right)^{\gamma},$ (4)
where $C=5.22\pm 0.61\,\mathrm{m\AA }$, $\beta=0.25\pm 0.10$, and $\gamma=0.31\pm 0.08$. This result further supports the connection between the cool CGM and the star formation activity of host galaxies.
@@
Figure 6: Equivalent width of AGN samples as a function of impact parameters. The shaded bands correspond to the two highest mass bins ($10^{9.9}$ and $10^{10.7}\,M_{\odot}$) from star-forming galaxies in Figure 3. The dotted line is calculated using Equation 3 with $M_{*}=10^{10.2}M_{\odot}$.
@@
Figure 8: Best-fit amplitudes as a function of median stellar masses for two redshift bins. The solid and open points represent low- and high-redshift samples, respectively. The solid and dashed lines in the left panel are the best-fit power laws for the low- and high-redshift samples.
Figure 9: Dependence of Ca II on the azimuthal angles for galaxies with $\log_{10}(M_{*}/M_{\odot})\geq 9.5$. Examples of the the galaxy–quasar pairs are shown at the top: from left to right are edge-on galaxies categorized into three azimuthal angles: disk, between disk and pole, pole; followed by face-on galaxies. The galaxies are located at the center, while the quasars are indicated by the cyan arrows. In the bottom panel, the gray band shows the measurements of the face-on samples.
@@ 3.2 Redshift dependence
With the correlations between Ca II absorption and galaxy properties explored, we now investigate the correlation between Ca II absorption and redshift as a function of stellar mass. We construct separate samples of star-forming and quiescent galaxies using redshift cuts at $z=0.15$ and $z=0.17$, respectively. In addition, stellar mass ranges are chosen so that redshift evolution has minimal impact across different mass bins. For star-forming galaxies, we include those with stellar masses in the range $10^{8.3}-10^{11.2}M_{\odot}$ at $z<0.15$, and $10^{9.4}-10^{11.2}M_{\odot}$ at $z>0.15$. For quiescent galaxies, we select those with stellar masses between $10^{9.5}$ and $10^{11.7}M_{\odot}$ at $z<0.17$, and between $10^{10.2}$ and $10^{11.7}M_{\odot}$ at $z>0.17$. The left and right panels of Figure 7 show the results for star-forming and quiescent galaxies, while the upper and lower panels show the results for the low- and high-redshift bins, respectively. We also obtain the best-fit amplitudes of power laws (Equation 2) with fixed $\alpha$ parameters obtained previously for star-forming and quiescent galaxies. The best-fit amplitudes are shown in Figure 8. The solid and open data points indicate the measurements at low and high redshifts, respectively.
@@
$A_{100}=C\left(\frac{M_{*}}{10^{10}\,M_{\odot}}\right)^{\beta}(1+z)^{\delta},$ (5)
and find $C=6.11\pm 1.58\,\mathrm{m\AA }$, $\beta=0.45\pm 0.10$, and $\delta=0.80\pm 1.57$. We note that despite the uncertainty, the redshift trend is similar to the redshift evolution observed in the cool gas around galaxies traced by Mg II absorption lines (e.g., Lan, 2020; Schroetter et al., 2021; Cherrey et al., 2025). For example, Lan (2020) showed that the covering fraction of Mg II absorption lines with $W_{\lambda 2796}>1\,\rm\AA$ around star-forming galaxies evolve with redshift as $(1+z)^{2.2\pm 0.4}$. On the other hand, we do not observe any redshift dependence for Ca II around quiescent galaxies, which is inconsistent with the results of Lan (2020) for quiescent galaxies. This might indicate that the cool gas around passive and star-forming galaxies evolve differently. However, further confirmation of such a scenario requires higher S/N measurements of Ca II absorption.
In Lan (2020), the author proposed that the redshift evolution of the Mg II covering fraction is connected with the redshift evolution of SFR of galaxies (e.g., Whitaker et al., 2012; Scoville et al., 2017; Koprowski et al., 2024) given that they have similar redshift evolution. For our Ca II measurements, we also estimate the SFR of star-forming galaxies at two redshift bins with stellar mass $10^{10}M_{\odot}$ and find that the ratio between the median SFR, $\mathrm{SFR}(z>0.15)/\mathrm{SFR}(z<0.15)$, is $\sim 1.6$. Together with our results indicating a SFR and absorption strength correlation with $\propto\rm SFR^{0.31}$ (Figure 5), the absorption strength difference ratio is expected to be $\sim 1.6^{0.31}\sim 1.16$, which is consistent with the redshift trend observed in Figure 8. This consistency suggests that the redshift evolution of Ca II absorption around star-forming galaxies might be closely related to the star formation activities of galaxies.
@@ 3.3 Azimuthal angle dependence
One can further investigate possible mechanisms contributing to gas observed in the CGM by measuring gas absorption around star-forming galaxies as a function of azimuthal angle. This is motivated by the fact that galaxy simulations predict that galactic outflows driven by supernova and black hole feedback preferentially escape galaxies along the minor axis (e.g., Nelson et al., 2019). To do so, we examine the azimuthal dependence of our sample by choosing star-forming galaxies with $\log_{10}\left({M_{*}/M_{\odot}}\right)\geq 9.5$ and with inclination angle larger than $60^{\circ}$. The inclination angles of galaxies and azimuthal angles between galaxy–quasar pairs are calculated based on the shape information measured via Tractor [^4] https://www.legacysurvey.org/dr9/catalogs/ (Lang et al., 2016) applied to the DESI Legacy Imaging Surveys (Dey et al., 2019). In general, the Tractor performs model-based photometry by simultaneously fitting sources across multiple bands and classifies them as point or extended sources with the point-spread function taken into account. Extended sources are further modeled using exponential, de Vaucouleurs, or composite profiles to best represent their light distribution. To evaluate the precision of the shape parameter measurements, we cross-match approximately 1500 DESI galaxies with inclination angles larger than $60^{\circ}$ to the Euclid Q1 data set (Euclid Collaboration et al., 2025), and compare the position angles used in this work and those from Euclid Q1. The result shows that approximately 95% of galaxies exhibit differences smaller than $10^{\circ}$. We then divide the selected sample into three azimuthal angle bins: $|\phi|\in[0^{\circ},30^{\circ}]$, $30^{\circ},60^{\circ}]$ and $60^{\circ},90^{\circ}]$, which correspond to the disk, between disk and pole, and the pole of the foreground galaxies. Galaxy image examples can be found in the upper panel of Figure 9, where the foreground galaxies are located at the centers, while the background quasars are pointed out by cyan arrows.
@@
The results are shown in the lower panel of Figure 9 with colors indicating measurements for different azimuthal angles and the gray band for the measurements of the face-on galaxies. There is no significant correlation between Ca II absorption and the azimuthal angle. This result is inconsistent with previous results of the cool CGM obtained at higher redshifts, showing preferentially stronger Mg II absorption along the minor axis than along the major axis within 50 kpc (e.g., Bordoloi et al., 2011; Bouché et al., 2012; Kacprzak et al., 2012; Lan et al., 2014; Lan & Mo, 2018; Zabl et al., 2019). In addition, our results are inconsistent with Ca II measurements from Zhu & Ménard (2013b). The authors reported that the absorption strength of Ca II is stronger along the minor axis than the major one. However, we note that the conclusion in Zhu & Ménard (2013b) is driven by a single measurement at $r_{p}\sim 15\ \mathrm{kpc}$, a trend that is not observed in our measurements with more pairs. We will further discuss the implications of our azimuthal angle measurements in Section 4.2.